OpenStudy (anonymous):

Factor This If you can show work 2x-18

OpenStudy (anonymous):

2x - 18 factor out the 2 2x / 2 = x and 18 / 2 = 9 so 2x-18 = 2(x-9) Answer: 2(x-9)